Issue Price

The following terms relate to independent bond issues:

  1. 400 bonds; $1,000 face value; 8% stated rate; 5 years; annual interest payments
  2. 400 bonds; $1,000 face value; 8% stated rate; 5 years; semiannual interest payments
  3. 820 bonds; $1,000 face value; 8% stated rate; 10 years; semiannual interest payments
  4. 2,150 bonds; $500 face value; 12% stated rate; 15 years; semiannual interest payments

Use the appropriate present value table:

PV of $1 and PV of Annuity of $1

Required:

Assuming the market rate of interest is 10%, calculate the selling price for each bond issue. If required, round your intermediate calculations and final answers to the nearest dollar.
Answer #1
Reqa.
Annual cash interest = 400,000*8% = 32,000
n = 5
I = 10%
Cashflows Amount PVF Present value
Annual cash interest 32,000 3.79079 121305.3
Principle outstanding 4,00,000 0.620921 248368.4
Price of the bonds 369673.7
Answer is $ 369,674
Req b.
Semi annual cash interest = 400,000*8%*6/12 = 16,000
n= 10
I = 5%
Cashflows Amount PVF Present value
Annual cash interest 16,000 7.72173 123547.7
Principle outstanding 4,00,000 0.613913 245565.2
Price of the bonds 369112.9
Answer is $ 369,113
Req c.
Semi annual cash interest = 400,000*8%*6/12 = 16,000
n= 20
I = 5%
Cashflows Amount PVF Present value
Annual cash interest 16,000 12.46221 199395.4
Principle outstanding 4,00,000 0.376889 150755.6
Price of the bonds 350151
Answer is $ 350,151
Req d.
Semi annual cash interest = 400,000*8%*6/12 = 16,000
n = 30
I = 5%
Cashflows Amount PVF Present value
Annual cash interest 16,000 15.37245 245959.2
Principle outstanding 4,00,000 0.231377 92550.8
Price of the bonds 338510
Answer is $ 338,510